OpenStudy (anonymous):

Simplify: (3x - 5) - (4x + 6) -x + 11 x + 11 -x -11 x - 11

OpenStudy (anonymous):

i got x+11, is that correct ?

OpenStudy (johnweldon1993):

Notquite...3x - 4x = -x and -5 - 6 = -11 so -x - 11 :)

OpenStudy (anonymous):

The first one

undeadknight26 (undeadknight26):

what he said.

OpenStudy (mathmale):

One of the easier ways of attacking this problem is to line up "like terms." In other words, we want all the x terms in one column and all the constant terms in another.

OpenStudy (anonymous):

thanks babes

OpenStudy (anonymous):

Late response :(

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (anonymous):

wait wait -x-11 or -x+11

OpenStudy (johnweldon1993):

Its -x - 11 the negative sign gets distributed through the second parenthesis

OpenStudy (mathmale):

(3x - 5) - (4x + 6), after changing the signs of the last two terms, becomes 3x - 5 -4x + 6 -------- It's not necessary to do the problem this way, but for some people it's a lot clearer.

OpenStudy (anonymous):

@PhantomX1 you said its the first one and the first one is -x+11

OpenStudy (anonymous):

@mathmale so whats the answer

OpenStudy (johnweldon1993):

Mathmale...look at your diagram again...you never changed the 6 to -6... promise...its -x - 11 lol :)

OpenStudy (mathmale):

I'm sure you won't have any problem adding those 2 expressions together yourself. Please combine: 3x-4x. then please combine -5+6

OpenStudy (mathmale):

@johnweldon1993 : Your eyes are sharper than mine. I have a prescription for new glasses and had better buy those glasses soon, don't you think? So, would you jump in, copy what I typed, fix it and post our shared new answer?

OpenStudy (anonymous):

@PhantomX1 hey, you're the one who said its -x+11

OpenStudy (anonymous):

Come on guyssss

OpenStudy (mathmale):

3x - 5 -4x + 6 <= wrong --------

OpenStudy (johnweldon1993):

Haha ehh quite good at math with or without it...just a sign change :) 3x - 5 -4x - 6 -------- -x - 11 <----this as I showed above...is your answer :)

OpenStudy (anonymous):

@mathmale can you just say what you think the answer is, dont complicate things

OpenStudy (mathmale):

Happiness is collaborating with others to obtain correct results. Right? :)

OpenStudy (anonymous):

@johnweldon1993 you were right by the way i got a 100, thanks

OpenStudy (mathmale):

If you say so. But wait 'til you're my age, and you might see things differently. Hope you're happy and satisfied with the outcome of your posting here.

OpenStudy (johnweldon1993):

Of course :) @yeah_i_suck_at_math

OpenStudy (anonymous):

@johnweldon1993 can you help me with other questions

OpenStudy (johnweldon1993):

sure @yeah_i_suck_at_math

OpenStudy (mathmale):

(3x - 5) - (4x + 6) = 3x-5-4x-6 = -4x+3x -5 -6 = -x - 11 This is just an expression; it's not an equation.

OpenStudy (anonymous):

Yea it is

OpenStudy (anonymous):

-x+11

OpenStudy (mathmale):

Always (well ... usually) welcome a good math discussion PhantomX: How about justifying your solution, while at the same time showing how and why you believe mine and johnweldon's is wrong?